OpenStudy (anonymous):

whats m-2n ------ 6 if m = -20 n= 5

OpenStudy (anonymous):

-------- means divide by right?

OpenStudy (anonymous):

no thats just a line

OpenStudy (anonymous):

 Like this?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

It means divided by. -20 - 2(5) = -20 - 10 = -30 Now divide this by 6 to get the answer.

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

You are right!!